# UVE: Are MLLMs Unified Evaluators for AI-Generated Videos?

### v2

@@ B.3 Adapting Single Ratings for Pairwise Comparison
The effectiveness of adapting single video ratings for video pair comparisons using InternVL-2.5 models is illustrated in Figure 9. The results aligns with our observations with Qwen2-VL (shown in Figure 5), revealing a model size-dependent pattern. While this adaptation strategy shows no benefit when applied to the 78B-scale model, it yields substantial improvements for the 8B-scale model.
B.4 Case Studies
@@
Table 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18 showcase UVE-Bench data examples of different subaspects, along with the evaluation results by InternVL-2.5-8B-MPO and InternVL-2.5-78B-MPO.
Model Model Size LLM Vision Encoder Frames per Video Video-LLaVA (Lin et al., 2023) 7B Vicuna-1.5 (Chiang et al., 2023) LanguageBind (Zhu et al., 2024) 8 LongVA-7B-DPO (Zhang et al., 2024a) 7B Qwen2 (Yang et al., 2024a) CLIP-ViT-L-336px (Radford et al., 2021) 16 ShareGPT4Video (Chen et al., 2024a) 8B LLaMA3 (Dubey et al., 2024) CLIP-ViT-L-336px (Radford et al., 2021) 16 VideoLLaMA2.1 (Cheng et al., 2024) 7B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 16 mPLUG-Owl3-7B (Ye et al., 2024) 7B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 16 VideoChat2-Mistral (Li et al., 2023b) 7B Mistral (Jiang et al., 2023) UMT-L/16 (Liu et al., 2022) 16 MiniCPM-V-2.6 (Yao et al., 2024) 7B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 16 LLaVA-OneVision-7B (Li et al., 2024) 7B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 12,16 LLaVA-OneVision-72B (Li et al., 2024) 72B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 12,16 LLaVA-Video-7B (Zhang et al., 2024b) 7B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 12,16 LLaVA-Video-72B (Zhang et al., 2024b) 72B Qwen2 (Yang et al., 2024a) SigLip-400M (Zhai et al., 2023) 12,16 Qwen2-VL-7B (Wang et al., 2024c) 7B - - 12,16 Qwen2-VL-72B (Wang et al., 2024c) 72B - - 12,16 InternVL-2.5-8B (Chen et al., 2024b) 8B InternLM2.5 (Cai et al., 2024) InternViT (Chen et al., 2023b) 12,16 InternVL-2.5-78B (Chen et al., 2024b) 78B InternLM2.5 (Cai et al., 2024) InternViT (Chen et al., 2023b) 12,16 GPT-4o-2024-08-06 (OpenAI, 2024) - - - 12,16
@@
Original Aspect Definition Evaluation Model UVE-Bench Aspect VideoScore Visual Quality the quality of the video in terms of clearness, resolution, brightness, and color VideoScore-v1.1 Overall Static Quality Aesthetic Quality Technical Quality Temporal Consistency the consistency of objects or humans in video VideoScore-v1.1 Appearance Consistency Dynamic Degree the degree of dynamic changes VideoScore-v1.1 Overall Dynamic Degree Text-to-Video Alignment the alignment between the text prompt and the video content VideoScore-v1.1 Overall Alignment Factual Consistency the consistency of the video content with the common-sense and factual knowledge VideoScore-v1.1 Structural Correctness Motion Naturalness VBench Overall Consistency overall video-text consistency ViCLIP (Wang et al., 2024d) Overall Alignment Motion Smoothness whether the motion in the generated video is smooth and follows the physical law of the real world AMT (Li et al., 2023d) Motion Naturalness Aesthetic Quality reflect aesthetic aspects such as the layout, the richness and harmony of colors, the photo-realism, naturalness, and artistic quality of the video frames LAION aesthetic predictor (LAION-AI, 2022) Aesthetic Quality Imaging Quality refers to the distortion (e.g., over-exposure, noise, blur) presented in the generated frames MUSIQ (Ke et al., 2021) Technical Quality Dynamic Degree the degree of dynamics (i.e., whether it contains large motions) RAFT (Teed & Deng, 2020) Overall Dynamic Degree Subject Consistency whether the subject appearance remains consistent throughout the whole video DINO (Caron et al., 2021) Appearance Consistency
@@
(a) Single Video Rating with InternVL2.5-8B-MPO.
(b) Video Pair Comparison with InternVL2.5-8B-MPO.
(c) Single Video Rating with Qwen2-VL-7B.
(d) Video Pair Comparison with Qwen2-VL-7B.
(e) Single Video Rating with Qwen2-VL-72B.
(f) Video Pair Comparison with Qwen2-VL-72B.
Figure 7: Results of different prompting strategies with InternVL-2.5-8B-MPO, Qwen2-VL-7B and Qwen2-VL-72B.
(a) Qwen2-VL-72B.
(b) InternVL-2.5-8B-MPO.
(c) InternVL-2.5-78B-MPO.
Figure 8: Results of different scoring strategies for single video rating with Qwen2-VL-72B, InternVL-2.5-8B-MPO and InternVL-2.5-78B-MPO.
@@
- Video 1: 0.076                       - Video 1: 0.085 - Video 2: 0.0028                     - Video 2: 0.047 - $\mathcal{A}^{\text{sinlge}}=1$                            - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: same bad        - Model Preference: Video 2 - Accuracy = ✓                          - Accuracy = ✗
@@
- Video 1: 1.2$\times$10-6                   - Video 1: 1.5$\times$10-3 - Video 2: 8.8$\times$10-8                   - Video 2: 3.4$\times$10-4 - $\mathcal{A}^{\text{sinlge}}=1$                              - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: same bad        - Model Preference: same bad - Accuracy = ✗                          - Accuracy = ✗
@@
- Video 1: 0.531                       - Video 1: 0.201 - Video 2: 0.004                       - Video 2: 0.047 - $\mathcal{A}^{\text{sinlge}}=1$                            - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 1        - Model Preference: Video 1 - Accuracy = ✓                        - Accuracy = ✓
@@
- Video 1: 0.500                       - Video 1: 0.651 - Video 2: 0.999                       - Video 2: 0.940 - $\mathcal{A}^{\text{sinlge}}=1$                            - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 2        - Model Preference: Video 2 - Accuracy = ✓                        - Accuracy = ✓
@@
- Video 1: 0.986                       - Video 1: 0.731 - Video 2: 0.991                       - Video 2: 0.755 - $\mathcal{A}^{\text{sinlge}}=1$                            - $\mathcal{A}^{\text{sinlge}}=0.319$ Video Pair Comparison
@@
- Model Preference: same good    - Model Preference: Video 2 - Accuracy = ✓                        - Accuracy = ✗
@@
- Video 1: 0.999                    - Video 1: 0.971 - Video 2: 0.971                    - Video 2: 0.915 - $\mathcal{A}^{\text{sinlge}}=8.6\times 10^{-6}$             - $\mathcal{A}^{\text{sinlge}}=1.9\times 10^{-5}$ Video Pair Comparison
@@
- Model Preference: Video 1        - Model Preference: Video 1 - Accuracy = ✗                        - Accuracy = ✗
@@
- Video 1: 0.990                    - Video 1: 0.971 - Video 2: 0.893                    - Video 2: 0.980 - $\mathcal{A}^{\text{sinlge}}=1$                         - $\mathcal{A}^{\text{sinlge}}=0$ Video Pair Comparison
@@
- Model Preference: Video 1        - Model Preference: Video 2 - Accuracy = ✓                       - Accuracy = ✗
@@
- Video 1: 0.852                    - Video 1: 0.731 - Video 2: 0.998                    - Video 2: 0.905 - $\mathcal{A}^{\text{sinlge}}=1$                         - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 2        - Model Preference: Video 2 - Accuracy = ✓                       - Accuracy = ✓
@@
- Video 1: 0.924                    - Video 1: 0.731 - Video 2: 0.986                    - Video 2: 0.798 - $\mathcal{A}^{\text{sinlge}}=0$                         - $\mathcal{A}^{\text{sinlge}}=0$ Video Pair Comparison
@@
- Model Preference: Video 2        - Model Preference: Video 2 - Accuracy = ✗                        - Accuracy = ✗
@@
- Video 1: 0.999                    - Video 1: 0.984 - Video 2: 0.009                    - Video 2: 0.755 - $\mathcal{A}^{\text{sinlge}}=1$                         - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 1        - Model Preference: Video 1 - Accuracy = ✓                       - Accuracy = ✓
@@
- Video 1: 0.269                    - Video 1: 0.076 - Video 2: 0.245                    - Video 2: 0.268 - $\mathcal{A}^{\text{sinlge}}=1$                         - $\mathcal{A}^{\text{sinlge}}=0$ Video Pair Comparison
@@
- Model Preference: same good    - Model Preference: Video 2 - Accuracy = ✗                        - Accuracy = ✗

### v4

@@ D.4 Adapting Single Ratings for Pairwise Comparison
The effectiveness of adapting single video ratings for video pair comparisons using InternVL-2.5 models is illustrated in Fig. 11. The results aligns with our observations with Qwen2-VL (shown in Fig. 4(d)), revealing a model size-dependent pattern. While this adaptation strategy shows no benefit when applied to the 78B-scale model, it yields substantial improvements for the 8B-scale model.
D.5 Effect of $\alpha$ and $\beta$
We have chosen default values of $\alpha=0.4$ and $\beta=0.8$ to represent the thresholds for categorizing videos as “bad” and “good”, respectively. To investigate the impact of these two parameters on single video rating evaluation, we conduct a sensitivity analysis. Specifically, we report the performance of four MLLMs across different values of $\alpha$ and $\beta$. As shown in Tab. 10 and Tab. 10, adjusting these thresholds reveals notable trends in performance: (1) Lowering $\beta$ (i.e., making the criteria for “good” more lenient) leads to increased agreement between MLLMs and humans in the “same good” category, resulting in higher performance. (2) Increasing $\alpha$ (i.e., making the criteria for “bad” stricter) also improves performance by enhancing alignment in the “same bad” category. However, despite the observed variations in performance as these parameters change, it is important to note that the relative performance of different MLLMs remains stable.
D.6 Case Studies
@@
Tab. 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22 showcase UVE-Bench data examples of different subaspects, along with the evaluation results by InternVL-2.5-8B-MPO and InternVL-2.5-78B-MPO.
@@
Model Model Size LLM Vision Encoder Frames per Video Video-LLaVA [34] 7B Vicuna-1.5 [11] LanguageBind [80] 8 LongVA-7B-DPO [72] 7B Qwen2 [65] CLIP-ViT-L-336px [47] 16 ShareGPT4Video [7] 8B LLaMA3 [12] CLIP-ViT-L-336px [47] 16 VideoLLaMA2.1 [10] 7B Qwen2 [65] SigLip-400M [71] 16 mPLUG-Owl3-7B [68] 7B Qwen2 [65] SigLip-400M [71] 16 VideoChat2-Mistral [31] 7B Mistral [22] UMT-L/16 [38] 16 MiniCPM-V-2.6 [67] 7B Qwen2 [65] SigLip-400M [71] 16 LLaVA-OneVision-7B [29] 7B Qwen2 [65] SigLip-400M [71] 12,16 LLaVA-OneVision-72B [29] 72B Qwen2 [65] SigLip-400M [71] 12,16 LLaVA-Video-7B [73] 7B Qwen2 [65] SigLip-400M [71] 12,16 LLaVA-Video-72B [73] 72B Qwen2 [65] SigLip-400M [71] 12,16 Qwen2-VL-7B [58] 7B - - 12,16 Qwen2-VL-72B [58] 72B - - 12,16 InternVL-2.5-8B [9] 8B InternLM2.5 [4] InternViT [8] 12,16 InternVL-2.5-78B [9] 78B InternLM2.5 [4] InternViT [8] 12,16 GPT-4o-2024-08-06 [45] - - - 12,16 Seed1.5-VL [15] 20B Act. Seed1.5-LLM Seed-ViT 12,16 Gemini2.5-Flash [14] - - - 12
@@
Original Aspect Definition Evaluation Model UVE-Bench Aspect VideoScore Visual Quality the quality of the video in terms of clearness, resolution, brightness, and color VideoScore-v1.1 Overall Static Quality Aesthetic Quality Technical Quality Temporal Consistency the consistency of objects or humans in video VideoScore-v1.1 Appearance Consistency Dynamic Degree the degree of dynamic changes VideoScore-v1.1 Overall Dynamic Degree Text-to-Video Alignment the alignment between the text prompt and the video content VideoScore-v1.1 Overall Alignment Factual Consistency the consistency of the video content with the common-sense and factual knowledge VideoScore-v1.1 Structural Correctness Motion Naturalness VBench Overall Consistency overall video-text consistency ViCLIP [59] Overall Alignment Motion Smoothness whether the motion in the generated video is smooth and follows the physical law of the real world AMT [33] Motion Naturalness Aesthetic Quality reflect aesthetic aspects such as the layout, the richness and harmony of colors, the photo-realism, naturalness, and artistic quality of the video frames LAION aesthetic predictor [28] Aesthetic Quality Imaging Quality refers to the distortion (e.g., over-exposure, noise, blur) presented in the generated frames MUSIQ [24] Technical Quality Dynamic Degree the degree of dynamics (i.e., whether it contains large motions) RAFT [55] Overall Dynamic Degree Subject Consistency whether the subject appearance remains consistent throughout the whole video DINO [5] Appearance Consistency
@@
Figure 5: Human baseline annotation interface.
(a) Results of different prompting strategies with InternVL-2.5-8B-MPO, Qwen2-VL-7B and Qwen2-VL-72B.
(b) Results of different scoring strategies for single video rating with Qwen2-VL-72B, InternVL-2.5-8B-MPO and InternVL-2.5-78B-MPO.
@@
Table 9: $\mathcal{A}^{single}$ results with varying values of $\beta$ when $\alpha$ is set to 0.4.
$\beta$=0.7 $\beta$=0.8 $\beta$=0.9
Qwen2-VL-2B 61.2 58.3 55.9
LongVA-DPO-7B 64.1 61.6 57.6
Qwen2-VL-7B 74.4 70.9 66.3
Qwen2-VL-72B 77.3 75.4 72.5
Table 10: $\mathcal{A}^{single}$ results with varying values of $\alpha$ when $\beta$ is set to 0.8.
$\alpha$=0.3 $\alpha$=0.4 $\alpha$=0.5
Qwen2-VL-2B 56.7 58.3 62.8
LongVA-DPO-7B 60.4 61.6 63.0
Qwen2-VL-7B 69.6 70.9 72.5
Qwen2-VL-72B 74.7 75.4 76.2
Table 11: Comparison of positive and negative prompts in terms of $\mathcal{A}^{single}$ with Qwen2-VL-7B as the backbone model.
Prompt Dynamic Degree Static Quality Temporal Quality Video-Text Alignment AVG Positive 84.6 64.6 51.1 85.4 70.9 Negative 66.0 62.7 55.0 16.9 50.9
@@
- Video 1: 0.076          - Video 1: 0.085
- Video 2: 0.0028         - Video 2: 0.047 - $\mathcal{A}^{\text{sinlge}}=1$            - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: same bad    - Model Preference: Video 2 - Accuracy =           - Accuracy =
@@
[…]
@@
- Model Preference: same bad    - Model Preference: same bad - Accuracy =           - Accuracy =
@@
- Video 1: 0.531          - Video 1: 0.201
- Video 2: 0.004          - Video 2: 0.047 - $\mathcal{A}^{\text{sinlge}}=1$            - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 1    - Model Preference: Video 1 - Accuracy =          - Accuracy =
@@
- Video 1: 0.500          - Video 1: 0.651
- Video 2: 0.999          - Video 2: 0.940 - $\mathcal{A}^{\text{sinlge}}=1$            - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 2    - Model Preference: Video 2 - Accuracy =          - Accuracy =
@@
- Video 1: 0.986          - Video 1: 0.731
- Video 2: 0.991          - Video 2: 0.755 - $\mathcal{A}^{\text{sinlge}}=1$            - $\mathcal{A}^{\text{sinlge}}=0.319$ Video Pair Comparison
@@
- Model Preference: same good    - Model Preference: Video 2 - Accuracy =          - Accuracy =
@@
- Video 1: 0.999          - Video 1: 0.971
- Video 2: 0.971          - Video 2: 0.915 - $\mathcal{A}^{\text{sinlge}}=8.6\times 10^{-6}$      - $\mathcal{A}^{\text{sinlge}}=1.9\times 10^{-5}$ Video Pair Comparison
@@
- Model Preference: Video 1     - Model Preference: Video 1 - Accuracy =          - Accuracy =
@@
- Video 1: 0.990          - Video 1: 0.971
- Video 2: 0.893          - Video 2: 0.980 - $\mathcal{A}^{\text{sinlge}}=1$           - $\mathcal{A}^{\text{sinlge}}=0$ Video Pair Comparison
@@
- Model Preference: Video 1     - Model Preference: Video 2 - Accuracy =          - Accuracy =
@@
- Video 1: 0.852          - Video 1: 0.731
- Video 2: 0.998          - Video 2: 0.905 - $\mathcal{A}^{\text{sinlge}}=1$           - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 2     - Model Preference: Video 2 - Accuracy =          - Accuracy =
@@
- Video 1: 0.924          - Video 1: 0.731
- Video 2: 0.986          - Video 2: 0.798 - $\mathcal{A}^{\text{sinlge}}=0$           - $\mathcal{A}^{\text{sinlge}}=0$ Video Pair Comparison
@@
- Model Preference: Video 2     - Model Preference: Video 2 - Accuracy =          - Accuracy =
@@
- Video 1: 0.999          - Video 1: 0.984
- Video 2: 0.009          - Video 2: 0.755 - $\mathcal{A}^{\text{sinlge}}=1$           - $\mathcal{A}^{\text{sinlge}}=1$ Video Pair Comparison
@@
- Model Preference: Video 1     - Model Preference: Video 1 - Accuracy =          - Accuracy =
@@
- Video 1: 0.269          - Video 1: 0.076
- Video 2: 0.245          - Video 2: 0.268 - $\mathcal{A}^{\text{sinlge}}=1$          - $\mathcal{A}^{\text{sinlge}}=0$ Video Pair Comparison
@@
- Model Preference: same good   - Model Preference: Video 2 - Accuracy =          - Accuracy =